OpenStudy (anonymous):

sin θ cos 3θ + cos θ sin 3θ = 0

OpenStudy (ranga):

Use the identity to simplify the left hand side: sin(A + B) = sin(A)cos(B) + cos(A)sin(B)

OpenStudy (jdoe0001):

http://www.cliffsnotes.com/assets/11025.jpg

OpenStudy (jdoe0001):

as ranga said, and then get the \(\bf sin^{-1}\) to both sides

OpenStudy (anonymous):

It says to simplify and then it says to find all the solutions. So when I simplify it I get sin4theta but im not sure what to do from there :X

OpenStudy (ranga):

sin θ cos 3θ + cos θ sin 3θ = 0 sin(θ + 3θ) = 0 sin(4θ) = 0 4θ = arcsin(0). Sine function is 0 when |θ| is 0, π, 2π, 3π, .... or when θ = ±n*π where n is an integer. 4θ = ±n*π θ = ±n * π/4 where n = 0, 1, 2, 3, 4,... (an integer)